LetX₁ ,X₂,X₃..., X₁₆ be a random sample from a N(μ, σ²) with σ² = 25 and sample mean X̅ = 60 . Find a 95% confidence interval for μ
1
Expert's answer
2019-03-27T14:36:31-0400

For a single mean from a normal distribution with known variance, a two-sided, 100(1 – α)% confidence interval is calculated by


"\bar{X}-z_{\alpha/2}*{\sigma \over \sqrt{n}}\le\mu\le\bar{X}+z_{\alpha/2}*{\sigma \over \sqrt{n}}"

For a 95 % confidence interval


"z_{\alpha/2}=z_{0.025}=1.96"

We have that


"\bar{X}=60, \sigma^2=25, n=16"


"60-1.96*{\sqrt{25}\over \sqrt{16}}\le\mu\le60+1.96*{\sqrt{25}\over \sqrt{16}}"

"57.55\le\mu\le62.45"

Hence


"95 \% CI \ [57.55, 62.45]"